OpenStudy (anonymous):

Write the sentence as a proportion. 3 diamonds is to 5 rubies as 12 diamonds is to 20 rubies

OpenStudy (anonymous):



OpenStudy (anonymous):

3/4 = 3*4/5*4 = 12/20